# Experimental neuromorphic computing based on quantum memristor

### v1

@@ III Experimental implementation
The main part of the experimental setup, fully represented in Fig. 2, is the physical quantum reservoir for information processing. It consists of an integrated waveguide circuit with 3 input/output modes and is designed according to a universal architecture [44]. The waveguides are fabricated by direct femtosecond laser writing [35, 45, 46] in an alumino-borosilicate glass substrate and feature single-mode operation at 1550 nm. Microstructured thermo-optic phase shifters are integrated on-chip, to provide efficient phase tunability [47]. Input and output ports are pigtailed with single-mode fibers.
The encoding of the classical input is set via an initial unitary transformation $\textbf{U}_{enc}$, through an on-chip MZI, as shown in Fig. 1b. For example, to encode the state $\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$, apply a phase $\phi=2\arccos(x)$ when injecting a photon into the lower mode of the MZI. However, since the PIC consists of three MZIs in total, we implement the product $\textbf{U}_{1}\cdot\textbf{U}_{enc}$ through the first MZI, as detailed in Fig. 2b.
The input single photon state is generated by a collinear Type-II spontaneous parametric down conversion source, pumped at 775 nm, which emits pairs of degenerate photons with a wavelength of 1550 nm. The employed non-linear crystal is a periodically poled titanyl phosphate, placed within a Sagnac interferometer. Since for this application we do not need entangled states, the crystal is pumped only in one spatial direction and generates the separable state $|01\rangle$. One photon is ancillary and heralds the generation of the second photon that is injected into the photonic processor. The three output modes of the photonic processor are connected to superconducting nanowire single photon detectors.
@@ IV Results
For the function prediction task three different cases are compared: (i) with quantum memristor, (ii) without memory (with no feedback loop) and (iii) with no reservoir. The prediction of $x^{n}$ with $n=4$ is shown in Fig. 3 and the inset plot shows the loss function of the readout unit for different $n$. Note that even without the feedback loop some non-linearity can be generated, due to the encoding from $x$ to $\rho_{x}$. However, the feedback loop enhances the impact of the non-linear measurement process and ultimately the performance of the algorithm. On the contrary, in the case of no reservoir, the output is simply a linear function, as the final ANN may only act linearly on the input.
@@
$L$ $y_{t}\approx\alpha u_{t}+\beta$ $2.76(29)$
$Q$ $y_{t}\approx\alpha u_{t}^{2}+\beta u_{t}+\gamma$ $2.05(22)$
$L+M$ $y_{t}\approx\alpha u_{t}+\beta u_{t-1}+\gamma$ $1.83(19)$
$Q+M$ $y_{t}\approx\alpha u_{t}^{2}+\beta u_{t-1}+\gamma$ $1.22(14)$
$L^{M}$ $y_{t}\approx\alpha u_{t}+\beta\;\text{Exp}_{4}(u_{t})+\gamma$ $1.14(11)$
$Q^{M}$ $y_{t}\approx\alpha u_{t}^{2}+\beta\;\text{Exp}_{4}(u_{t})+\gamma$ $0.78(8)$
@@
Table 1: Mean Squared Errors for classical machine learning models. This table shows the average MSE achieved by several classical models, in comparison to the one presented in this work, indicated by QMEM. The numbers in the bracket indicate one standard deviation of 100 runs. In detail, $L$ and $Q$ stand, respectively, for models that can base their predictions on linear and quadratic manipulation of the input. $L+M$ ($Q+M$) indicates the same models, but with access to the previous element of the output series. In $L^{M}$ ($Q^{M}$), instead, the access to previous elements is extended to the exponential moving average of all the past outputs, with a memory extent of 4. The reported values are obtained through numerical simulations.
In Fig. 4, we show the experimental prediction of the network for the previously described NARMA task. For the case with no feedback loop, each output depends only on the last input, since only the feedback loop allows storing the information of previous data points. It is noteworthy that the case with no feedback loop still has decent performance, due to the low non-linearity of the task (see section IC of the Supplemental Material for further details). To have a more complete overview of the performance of our algorithm, we compared it with similar classical algorithms. In particular, we tested the aforementioned tasks with models that could base their predictions on a linear or quadratic manipulation of the inputs in three cases: with no memory, with access to the previous outcome and to the exponential moving average of all past inputs, with memory extent equal to 4. For all these cases, the average MSE obtained in our case is lower, highlighting the enhancement brought by the quantum memristor (the numerical values are reported in Table I).
@@ V Discussion
In this work, we have implemented the first instance of QRC exploiting single photon states based on a quantum memristor device. These machine learning models work by injecting input data into a fixed random reservoir, which is then followed by a one-layered linear ANN (the only part of the model which is trained) [34]. The required non-linearity is provided by the evolution of the quantum system through the photonic quantum memristor.
The last part of the model, featuring an ANN, is implemented through standard computing. We tested our model on two function prediction tasks: several power functions and the NARMA task. Due to its regularity, the first case is only to test the non-linearity that is achievable through our apparatus. The latter instead tests also the memory of our model, given that its outcomes depend on previous ones.
@@
Ours is the first example of neuromorphic architectures exploiting single-photon states (encoded in the path degree of freedom) and implementing a non-linear behaviour through an adaptive protocol, performed by the quantum memristor. Moreover, we carry out a systematic analysis to distinguish the source of such a non-linearity, i.e. the encoding or the feedback loop. For this reason, for all of the tasks, we compare the case featuring the quantum memristor (QRC), to the case without feedback loop. For all of the cases, the quantum memristor brings an enhanced accuracy with respect to classical models. This scheme, although applied to relatively simple tasks, can be seen as a building block for more complex networks, as the output of the memristor, tracing out the update mode, preserves coherence and can be used as an input for successive layers (see section II of the Supplemental Material for further details).
This study paves the way for further investigations related to the non-linearities achievable on photonic platforms. This is particularly relevant, considering that hybrid optical-electronic ANNs have been proven to require a lower energy consumption than standard ones [25, 26].
@@
In this framework, the memristor could be employed to constitute the activation function layer of quantum neural networks. Another interesting feature is that a quantum reservoir could be used to directly process quantum systems, tackling tasks with no classical equivalent. For instance, to investigate foundational aspects or quantum state properties [49, 50, 51]. Further studies can involve the implementation of a quantum feedback, where previous states interfere with new ones, through a loop architecture [52]. Analogously, the implementation of a fast feedback would allow to make coherent adjustments to the quantum state. This would allow for a larger memory storage and open up to the possibility of tackling a wider range of tasks.
@@ Acknowledgments
The fabrication of the photonic integrated circuit was partially carried out at PoliFAB (https://www.polifab.polimi.it/), the micro- and nano-fabrication facility of Politecnico di Milano. F.C. and R.O. would like to thank the PoliFAB staff for technical advice and support. This research was funded in whole or in part by the Austrian Science Fund (FWF)[10.55776/ESP205] (PREQUrSOR), [10.55776/F71] (BeyondC), [10.55776/FG5] (Research Group 5) and [10.55776/I6002] (PhoMemtor).
@@
The authors declare no competing interest.

### v2

@@ III Experimental implementation
The main part of the experimental setup, fully represented in Fig. 2, is the physical quantum reservoir for information processing. It consists of an integrated waveguide circuit with 3 input/output modes and is designed according to a universal architecture [48]. The waveguides are fabricated by direct femtosecond laser writing [32, 49, 50] in an alumino-borosilicate glass substrate and feature single-mode operation at 1550 nm. Microstructured thermo-optic phase shifters are integrated on-chip, to provide efficient phase tunability [51]. Let us highlight that, due to the limited size, our circuit does not suffer from thermal crosstalk (for more detailed information about the calibration procedure, see the Supplementary Material note III). Furthermore, due to the low birefringence of the substrate, this circuit is polarization independent. Input and output ports are pigtailed with single-mode fibers.
The encoding of the classical input is set via an initial unitary transformation $\textbf{U}_{enc}$, through an on-chip MZI, as shown in Fig. 1b. For example, to encode the state $\sqrt{x}|0\rangle+\sqrt{1-x}|1\rangle$, apply a phase $\phi=2\arccos(x)$ when injecting a photon into the lower mode of the MZI. However, since the optical circuit consists of three MZIs in total, we implement the product $\textbf{U}_{1}\cdot\textbf{U}_{enc}$ through the first MZI, as detailed in Fig. 2b.
The input single photon state is generated by a collinear Type-II spontaneous parametric down conversion source, pumped at 775 nm, which emits pairs of degenerate photons with a wavelength of 1550 nm. The employed nonlinear crystal is a periodically poled titanyl phosphate, placed within a Sagnac interferometer. Since for this application we do not need entangled states, the crystal is pumped only in one spatial direction and generates the separable state $|01\rangle$. One photon is ancillary and heralds the generation of the second photon that is injected into the photonic processor. This implies that we post-select our statistics, considering only the cases where both the heralding and signal photons are detected. The three output modes of the photonic processor are connected to superconducting nanowire single photon detectors.
@@ IV Results
For the monomial prediction tasks three different cases are compared: (i) with quantum memristor, (ii) without memory (with no feedback loop) and (iii) with no reservoir. The prediction of $x^{n}$ with $n=4$ is shown in Fig. 3 and the inset plot shows the loss function of the readout unit for different $n$. Note that even without the feedback loop some nonlinearity can be generated, due to the encoding from $x$ to $\rho_{x}$. However, the feedback loop enhances the impact of the nonlinear measurement process and ultimately the performance of the algorithm. Let us highlight that, when comparing the cases with and without feedback, we are picking the best hyperparameters (namely $U_{1}$ and $U_{3}$ and the coefficients of the feedback function), separately for the two cases, to have a fair comparison. In contrast, in the case of no reservoir, the output is simply a linear function, since the final linear regression only gives a weighted sum of the inputs.
@@
$L$ $y_{t}\approx\text{Poly}_{1}(x_{t})$ $2.76(29)$
$C$ $y_{t}\approx\text{Poly}_{3}(x_{t})$ $2.05(22)$
$L+M$ $y_{t}\approx\text{Poly}_{1}(x_{t},x_{t-1})$ $1.83(19)$
$C+M$ $y_{t}\approx\text{Poly}_{3}(x_{t},x_{t-1})$ $0.92(11)$
@@
Table 1: Mean Squared Errors for classical machine learning models for NARMA task. This table shows the average MSE achieved by several classical models, which have access to a similar amount of resources, in terms of memory and free variables, in comparison to the one presented in this work, indicated by QMEM. The numbers in the bracket indicate one standard deviation of 100 runs. $\text{Poly}_{d}(a_{1},\ldots,a_{n})$ is the polynomial of variables $\{a_{1},\ldots,a_{n}\}$ of degree at most $d$. In detail, $L$ and $C$ stand, respectively, for models that can base their predictions on linear and cubic manipulation of the input. The first features 2 and the second 4 free variables. $L+M$ ($C+M$) indicates the same models, but with access to the previous element of the output series. In this case, the first features 3 and the second 9 free variables. For an accurate description of such models, see Supplementary Material note II. The reported values are obtained through numerical simulations. We note that polynomials of higher order improve neither of the results. Our model features only 1 free variable, given by the memory decay $m_{d}$ (see Eq. (4)).
In Fig. 4, we show the experimental prediction of the network for the previously described time series prediction tasks. For the case with no feedback loop, each output depends only on the last input, since there is no storing the information of previous data points. It is worth noting that cases without feedback loops still achieve fairly good results (although worse than the case where the quantum memristor is employed). For the NARMA task, this is due to its low nonlinearity, while for the Mackey-Glass and Santa Fe tasks, it results from the high density of probing points, which makes the prediction almost linear. To have a more complete overview of the performance of our algorithm, we present the results achieved on the same tasks by simple classical algorithms. First, we consider a model without memory, where the prediction is a polynomial function of the input data. In particular, we consider two models, featuring 2 and 4 free parameters. Next, we examine two models with the simplest possible memory, which records only the previous input. Hence, the final prediction will be given by a polynomial function of the current and previous input. In this case, the two models have, respectively, 3 and 9 free parameters. Our model featuring the quantum memristor has a slightly better performance than those models (see Table 1), even when the classical one has access to past inputs. For the comparison of the other tasks and the mathematical description of these models, see Supplementary Material note IIC.
@@ V Discussion
In this work, we have implemented the first instance of QRC exploiting single photon states based on a quantum memristor device. These machine learning models work by injecting input data into a fixed random quantum reservoir, implemented through a photonic quantum memristor, which is then followed by a classical linear regression model [31]. We tested our algorithm on four function prediction tasks: several monomial functions, and then NARMA, Santa Fe and Mackey-Glass time series. Due to its regularity, the first case is only to test the nonlinearity that is achievable through our apparatus. The others, instead, test also the memory of our model, given that its outcomes depend on previous ones.
@@
Ours is the first example of neuromorphic architectures exploiting single-photon states (encoded in the path degree of freedom) and implementing a nonlinear behaviour through an adaptive protocol, performed by the quantum memristor. Moreover, we carry out a systematic analysis to distinguish the source of such a nonlinearity, i.e. the encoding or the feedback loop. For this reason, for all of the tasks, we compare the case featuring the quantum memristor (QRC), to the case without feedback loop, consistently showing that the quantum memristor brings an enhanced accuracy with respect to classical models. This scheme, although applied to relatively simple tasks, can be seen as a building block for more complex networks, as the output of the memristor, when tracing out the update mode, preserves coherence and can be used as an input for successive layers (see the Supplementary Material note I for further details).
@@
A potential way of enhancing the expressivity of the proposed model is to use multiple memristors in a consecutive way, as shown for instance in [53], and updating both the internal variables of those memristors, as well as the input state, so that new inputs would interfere with states encoding previous outcomes. Analogously, the implementation of a fast feedback would allow to make coherent adjustments to the quantum state. This would result in a larger memory storage and open up to the possibility of tackling a wider range of tasks. Moreover, feeding parts of the outputs as new inputs, for instance through a loop architecture [54], could unlock the possibility of tackling forecasting tasks [55]. Another interesting feature is that a quantum reservoir could be used to directly process quantum systems, tackling tasks with no classical equivalent. For instance, to investigate foundational aspects or quantum state properties [29, 56, 57].
Besides exploring the possibilities unlocked by the quantum features of the memristor, this study paves the way for further investigations related to the nonlinearities achievable on photonic platforms, for optical computing. This is particularly relevant, considering that hybrid optical-electronic ANNs have been proven to require a lower energy consumption than standard ones [25, 26]. In this framework, the memristor could be employed to constitute the activation function layer of quantum neural networks and, after further investigations on the relation between the memristor’s dynamics with neurons one, it could unlock the possibility of implementing spiking optical neural networks [58, 59, 60].
@@ Acknowledgments
The fabrication of the photonic integrated circuit was partially carried out at PoliFAB (https://www.polifab.polimi.it/), the micro- and nano-fabrication facility of Politecnico di Milano. F.C. and R.O. would like to thank the PoliFAB staff for technical advice and support. This research was funded in whole or in part by the Austrian Science Fund (FWF)[10.55776/ESP205] (PREQUrSOR), [10.55776/F71] (BeyondC), [10.55776/FG5] (Research Group 5) and [10.55776/I6002] (PhoMemtor). R.O. acknowledges financial support by ICSC – National Research Center in High Performance Computing, Big Data and Quantum Computing, funded by European Union– NextGenerationEU.
Author Contribution
Mir.S. and I.A. designed and conducted the experiment, Mich.S. developed the theory and algorithm for the time series predictions and analyzed the corresponding data, I.A. developed the algorithm for power predictions and I.A. and Mir.S. analyzed the corresponding data. R.A., A.C. and F.C. conducted the design, fabrication and calibration of the integrated photonic processor. Mir.S., I.A. and Mich.S. wrote the first draft of the manuscript. I.A., Mag.S., R.O., B.D. and P.W. supervised the whole project. All authors discussed the results and reviewed the manuscript.
@@
F.C. and R.O. are co-founders of the company Ephos. The authors declare that they have no other competing interests.